# arXiv:2603.01338v1[math.AP] 02 Mar 2026

### v1

@@ 3. Bilinear Dispersive Estimates
$\lesssim$ $\biggl\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)f](\xi-\eta)$
$\times{{\mathcal{F}}}\left[\frac{\partial_{x_{1}}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})\{3\partial_{x_{1}}^{4}+6(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})\partial_{x_{1}}^{2}+(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})^{2}\}}{\Delta^{2}(3\partial_{x_{1}}^{2}-\partial_{x_{2}}^{2}-\partial_{x_{3}}^{2})^{3}}V(t)g\right](\eta)d\eta d\tau\biggl\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)x_{1}f](\xi-\eta){{\mathcal{F}}}\left[\frac{\partial_{x_{1}}^{2}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})}{p(-i\nabla)}V(t)g\right](\eta)d\eta d\tau\right\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)f](\xi-\eta){{\mathcal{F}}}\left[\frac{\partial_{x_{1}}^{2}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})}{p(-i\nabla)}V(t)x_{1}g\right](\eta)d\eta d\tau\right\|_{L_{\xi}^{2}}.$
By the decay estimate (Lemma 2.1 (2.2)) and the boundedness of $\Delta^{-2}\{3\partial_{x_{1}}^{4}+6(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})\partial_{x_{1}}^{2}+(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})^{2}\}$ and $\Delta^{-1}(\partial_{x_{2}}^{2}+\partial_{x_{3}}^{2})$ in $L^{2}$, we have
@@
where the norm $\|\cdot\|_{X_{\delta}}$ is given by (1.5).
@@
We next derive an a priori estimates for $w_{\lambda,\mu}$ independent of $\lambda$ and $\mu$ under the assumption that $r:=\|u_{+}\|_{X_{\delta}}<\infty$ with $0<\delta<1$. We abbreviate $w_{\lambda,\mu}$ to $w$. Let
@@
Hence, from (4), (4) and (4), we see that for any $M>0$, we obtain
@@
$+t^{-3\alpha+\frac{\nu}{3}+\frac{1}{2}}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{2}\left(\sup_{t\in[T,\infty)}t^{\alpha}\||\partial_{x_{1}}|^{\frac{\nu}{2}}w(\tau)\|_{L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{2}{\nu}})}\right)$
@@
$\|w(t)\|_{H_{x}^{3}}^{2}\sim\|w(t)\|_{H_{x}^{3}}^{2}+6\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx+M\|w\|_{L_{x}^{2}}^{14}.$
Hence we have (4.1). ∎

### v2

@@ 3. Bilinear Dispersive Estimates
$\lesssim$ $\biggl\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)f](\xi-\eta){{\mathcal{F}}}\left[(\partial_{\eta_{1}}B_{1,1})(-i\nabla)V(t)g\right](\eta)d\eta d\tau\biggl\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)x_{1}f](\xi-\eta){{\mathcal{F}}}\left[B_{1,1}(-i\nabla)V(t)g\right](\eta)d\eta d\tau\right\|_{L_{\xi}^{2}}$
$+\left\|\xi_{1}\int_{t}^{\infty}\!\!\!\int_{\mathbb{R}^{3}}\tau^{-1}{{\mathcal{F}}}[V(t)f](\xi-\eta){{\mathcal{F}}}\left[B_{1,1}(-i\nabla)V(t)x_{1}g\right](\eta)d\eta d\tau\right\|_{L_{\xi}^{2}}.$
By the decay estimate (Lemma 2.1 (2.2)) and the inequality for $B_{1,1}$ in Lemma 3.2, we have
@@
where the semi-norm $\|\cdot\|_{X_{\delta}}$ is given by (1.5).
@@
We next derive an a priori estimates for $w_{\lambda,\mu}$ independent of $\lambda$ and $\mu$ under the assumption that $r:=\|u_{+}\|_{X_{\delta}}<\infty$, where $\|\cdot\|_{X_{\delta}}$ is defined by (1.5). We abbreviate $w_{\lambda,\mu}$ to $w$. Let
@@
Hence, from (4), (4) and (4), we see that for any $M>0$,
@@
$+t^{-3\alpha+\frac{\nu}{3}+\frac{1}{2}}\left(\sup_{t\in[T,\infty)}t^{\alpha}\|w(t)\|_{H_{x}^{3}}\right)^{2}$
$\qquad\qquad\qquad\quad\quad\times\left(\sup_{t\in[T,\infty)}t^{\alpha}\||\partial_{x_{1}}|^{\frac{\nu}{2}}w(\tau)\|_{L_{\tau}^{\frac{6}{3-2\nu}}(t,\infty;W_{x}^{2,\frac{2}{\nu}})}\right)$
@@
(4.14) $\|w(t)\|_{H_{x}^{3}}^{2}\sim\|w(t)\|_{H_{x}^{3}}^{2}+6\int_{\mathbb{R}^{3}}w(\Delta w)^{2}dx+M\|w\|_{L_{x}^{2}}^{14}.$
Combining (LABEL:y1) with (4.14), we have (4.1). ∎
@@
Appendix A Proof of Lemma 3.2.
In this appendix, we prove Lemma 3.2.
Proof of Lemma 3.2..
By (3.3), we see
(A.1) $\partial_{\eta_{1}}\phi(\xi,\eta)$ $=$ $3\xi_{1}^{2}+\xi_{2}^{2}+\xi_{3}^{2}-6\xi_{1}\eta_{1}-2\xi_{2}\eta_{2}-2\xi_{3}\eta_{3},$
(A.2) $\partial_{\eta_{2}}\phi(\xi,\eta)$ $=$ $2\xi_{1}\xi_{2}-2\xi_{1}\eta_{2}-2\eta_{1}\xi_{2},$
(A.3) $\partial_{\eta_{3}}\phi(\xi,\eta)$ $=$ $2\xi_{1}\xi_{3}-2\xi_{1}\eta_{3}-2\eta_{1}\xi_{3}.$
Hence, we have
Therefore,
(A.4) $\eta_{2}\xi_{2}+\eta_{3}\xi_{3}=-\frac{3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}}{2\eta_{1}}\xi_{1}+N_{1},$
where $N_{1}=N_{1}(\xi,\eta)$ is given by
(A.5) $N_{1}=\frac{1}{2\eta_{1}}\left(\phi(\xi,\eta)-\eta\cdot\nabla_{\eta}\phi(\xi,\eta)\right).$
[…]
Therefore,
(A.6) $\eta_{3}\xi_{2}-\eta_{2}\xi_{3}$ $=$ $\frac{1}{2\eta_{1}}\left\{(\xi_{3}-\eta_{3})\partial_{\eta_{2}}\phi-(\xi_{2}-\eta_{2})\partial_{\eta_{3}}\phi\right\}.$
By (A.4) and (A.6),
where $N_{2}=N_{2}(\xi,\eta)$ is given by
Hence
(A.8) $\xi_{2}^{2}+\xi_{3}^{2}=\frac{(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})^{2}}{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})}\xi_{1}^{2}+N_{2}(\xi,\eta).$
Substituting (A.4) and (A.8) into (A.1), we have
$\partial_{\eta_{1}}\phi(\xi,\eta)$
Therefore
$=$ $\partial_{\eta_{1}}\phi+2N_{1}-N_{2}.$
[…]
Hence by (A.4), we obtain
Therefore,
By ${(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})/\eta_{1}^{2}\times}$ $\{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}\}/\{4\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})\}\times(A)$, we have
where $N_{3}=N_{3}(\xi,\eta)$ is given by
Hence we obtain
(A.12) $\xi_{1}=\frac{4\eta_{1}^{3}(\eta_{2}^{2}+\eta_{3}^{2})}{3(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{2}}N_{3}.$
Substituting (A.5), (LABEL:n2), (LABEL:n3) into (A.12), we have (3.5), where
$A_{0}(\eta)$ $=$ $\frac{1}{3p(\eta)}\{9\eta_{1}^{4}+30\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+5(\eta_{2}^{2}+\eta_{3}^{2})^{2}\},$
$A_{1}(\eta)$ $=$ $-\frac{1}{3p(\eta)}\eta_{1}\{9\eta_{1}^{4}+18\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2})+(\eta_{2}^{2}+\eta_{3}^{2})^{2}\},$
$A_{j}(\eta)$ $=$ $-\frac{4}{3p(\eta)}\eta_{i}(\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
$B_{0,1}(\eta)$ $=$ $-\frac{4}{p(\eta)}\eta_{1}(\eta_{2}^{2}+\eta_{3}^{2}),$
$B_{0,k}(\eta)$ $=$ $-\frac{2}{3p(\eta)}\eta_{k}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad k=2,3,$
$B_{1,1}(\eta)$ $=$ $\frac{4}{p(\eta)}\eta_{1}^{2}(\eta_{2}^{2}+\eta_{3}^{2}),$
$B_{1,k}(\eta)$ $=$ $\frac{2}{3p(\eta)}\eta_{1}\eta_{k}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad k=2,3,$
$B_{j,1}(\eta)$ $=$ $\frac{4}{p(\eta)}\eta_{1}\eta_{j}(\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
$B_{j,2}(\eta)$ $=$ $\frac{2}{3p(\eta)}(\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
$C_{j,1}(\eta)$ $=$ $-\frac{2}{3p(\eta)}\eta_{5-j}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
$C_{j,2}(\eta)$ $=$ $\frac{2}{3p(\eta)}\eta_{j}(3\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2}),\quad j=2,3,$
with $p(\eta)=(\eta_{1}^{2}+\eta_{2}^{2}+\eta_{3}^{2})(3\eta_{1}^{2}-\eta_{2}^{2}-\eta_{3}^{2})^{2}$. Hence we have Lemma 3.2. ∎